OpenStudy (anonymous):

Did I simplify this Algebra problem correctly? 2g^6t^-6 this is 2gt^-36, correct?

hartnn (hartnn):

is it \[2g^6t^{-6}\]?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

use this: \[t^{-6}=\frac{1}{t^6}\] and then \[\frac{g^6}{t^6}=(\frac{g}{t})^6\]

OpenStudy (anonymous):

Sorry, my computer froze.

OpenStudy (anonymous):



hartnn (hartnn):

ehy did 2 go int he denominator?it stays in numerator....

OpenStudy (anonymous):